OpenStudy (anonymous):

write the standard form of the equation of the circle (-9,7) and a radius of 10

OpenStudy (anonymous):

(x-(-9))^2+(y-7)^2=(10)^2 (x+9)^2+(y-7)^2=100

OpenStudy (anonymous):

\[(x+9)^2+(y-7)^2=100\]